OpenStudy (anonymous):

A coin is tossed three times. Find the probability of getting 2 heads

OpenStudy (amistre64):

P(hht) or P(thh) or P(hth)

OpenStudy (amistre64):

since the probability of getting h is 1/2 for any toss the probability of getting any 3 results is .5^3 3 ways to get 2h, so 3*.5^3 seems right to me

OpenStudy (anonymous):

uh haha the answer is 3/8 btw

OpenStudy (anonymous):

how is that the answer?

OpenStudy (amistre64):

yes :) 3*(1/2)^3 = 3/2^3 = 3/8

OpenStudy (amistre64):

well, lets table it up that might help to see it happen instead of theory based

OpenStudy (amistre64):

list all the possible ways of getting results from 3 tosses hhh hht hth thh htt tht tth ttt so we have 8 possible outcomes, 3 of which have only 2 heads in them 3 out of 8 is 3/8

OpenStudy (anonymous):

isn't there 9 possible outcomes?

OpenStudy (anonymous):

wait nvm LOL

OpenStudy (anonymous):

ok I get it thx

OpenStudy (amistre64):

did i miss one or did you miscount?

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thanks